# Secure Multi-Key Homomorphic Encryption with Application to Privacy-Preserving Federated Learning

### v1

@@ VI-C2 Experimental Results
Training to Convergence. We next compare the total training time and network traffic of THE-, CDKS-, and SMHE-based PPFL models until convergence is reached. The number of clients is set to $10$. Due to the exceedingly high time and communication costs involved in the full training process, we simulate federated learning locally until convergence and estimate the total training time and communication cost based on the bandwidth and iteration counts. Table III presents the estimated results for both FCN and AlexNet models, from which we can observe that (1) All PPFL models using SMHE, CDKS, and THE achieve high prediction accuracy, exceeding $97\%$ on FCN and around $74\%$ on AlexNet. (2) For the FCN model, SMHE achieves comparable accuracy ($97.90\%$) to the plaintext baseline ($97.94\%$), with only slightly more training epochs ($35$ vs. $34$). It incurs $6.22$ hours of training time and $41.40$ GB network traffic, which are less than 2×2$\times$2 × that of THE and CDKS. (3) For the more complex AlexNet model, SMHE also reaches similar accuracy to the plaintext case ($73.84\%$ vs. $74.01\%$) with slightly fewer epochs. However, it requires $232.49$ hours and $1397.31$ GB of traffic, which are less than 1.8×1.8$\times$1.8 × of THE and CDKS. Overall, compared with THE and CDKS, the SMHE-based PPFL models introduce only a modest increase in training time and communication cost (less than 2×2$\times$2 ×) while ensuring stronger security guarantees and offering better dynamism, making it a practical and secure choice for PPFL applications.
VII Conclusion
In this paper, we propose a novel secure multi-key homomorphic encryption (SMHE) scheme tailored for privacy-preserving FL. Our scheme addresses critical security vulnerabilities identified in previous MKHE schemes, specifically those proposed by Chen et al. and Kim et al., which inadvertently reveal plaintext information during multiparty secure computation tasks such as federated learning. By introducing a masking scheme into the CKKS and BFV frameworks, our enhanced scheme ensures the confidentiality of data while supporting efficient and secure homomorphic operations. Future work includes further optimizing the efficiency of ciphertext expansion in SMHE and exploring its integration with advanced applications such as secure model inference and encrypted machine learning.

### v2

@@ 6.5.2 Experimental Results
Training to Convergence. We next compare the total training time and network traffic of different PPFL models until convergence is reached. The number of clients is set to $10$. Due to the exceedingly high time and communication costs involved in the full training process, we simulate federated learning locally until convergence and estimate the total training time and communication cost based on the bandwidth and iteration counts. Table III presents the estimated results for both FCN and AlexNet models, from which we can observe that (1) All PPFL models using SMHE, CDKS, and THE achieve high prediction accuracy, exceeding $97\%$ on FCN and around $74\%$ on AlexNet. (2) For both the FCN and AlexNet models, SMHE achieves comparable accuracy to the plaintext baseline ($97.90\%$ vs. $97.94\%$ for FCN and $73.84\%$ vs. $74.01\%$ for AlexNet) and it incurs 6.226.22/$232.49$ hours of training time and 41.4041.40/$1397.31$ GB network traffic for FCN/AlexNet, which are less than $2\times$ that of THE and CDKS. Overall, compared with THE and CDKS, the SMHE-based PPFL models introduce only a modest increase in training time and communication cost (less than $2\times$) while ensuring stronger security guarantees and offering better dynamism, making it a practical and secure choice for PPFL applications.
7 Conclusion
In this paper, we propose a novel secure multi-key homomorphic encryption (SMHE) scheme. Our scheme addresses critical security vulnerabilities identified in previous MKHE schemes, specifically those proposed by Chen et al. and Kim et al., which inadvertently reveal plaintext information during multiparty secure computation tasks such as federated learning. By introducing a masking scheme into the CKKS and BFV schemes, our enhanced scheme ensures the confidentiality of data while supporting efficient and secure homomorphic operations. Future work includes further optimizing the efficiency of ciphertext expansion in SMHE and exploring its integration with advanced applications such as secure model inference and encrypted machine learning.